OpenStudy (johnnydicamillo):

Find the volume of the solid obtained by rotating the region bounded by the given curves about the specified axis

OpenStudy (johnnydicamillo):

x = 4y^2 y = 1, x =0. I know most of the formulas, I just need some help getting started

OpenStudy (johnnydicamillo):

for example I am given three different lines to graph, and do I need to get convert anything first from x to y or y to x? and which ones do I use to find the upper and lower limits?

OpenStudy (johnnydicamillo):

I know the formula\[\int\limits_{a}^{b}\pi [(outer)^2 - (inner)^2]\]

OpenStudy (anonymous):

What is the "specified axis?"

OpenStudy (johnnydicamillo):

the y axis

OpenStudy (anonymous):

Depending on the method of integration you wish to use, you will be integrating with respect to either \(x\) or \(y\). Either method works, so it's up to you. It looks like you have the disk method in mind?